Put the functions below in the rate of growth.

1. 100·3x

2. x!/1000

3. x4 · log(x)

4. x5

5. 200log(x)
Answer #1
Order from low to high order of growth
200log(x) = O(log(x))
100·3x = O(x)
x4 · log(x) = O(x4 · log(x)) 
x5 = O(x5)
x!/1000 = O(x!)